The speed of an ultrasonic sound of frequency 54kHz in air is 345.81m/s. What is the air temperature?
Answer #1

We know,

v_{sound}=331.4+0.6T

so, 345.81 = 331.4+0.6T

or 0.6 T = 345.81-331.4

or 0.6 T = 14.41

T = 14.41/0.6 = 24.017 C